OpenStudy (anonymous):

Please help! Simplify -(x-4) ----- 4-x

OpenStudy (anonymous):

equals 1

OpenStudy (anonymous):

\[\frac {-(x-4)} {4-x} = \frac {4-x} {4-x} = 1 \]

OpenStudy (anonymous):

Thank you for the explanation

OpenStudy (anonymous):

You're most welcome